OpenStudy (littlebird):

In the position formula x(t)=At2 + Bt + C, what is the value of the acceleration in terms of A?

OpenStudy (littlebird):

Am I solving for A? Deriving?

OpenStudy (kittiwitti1):

Is this calculus level?

OpenStudy (solomonzelman):

If this is calculus, then the acceleration is the second derivative of the position with respect to time.

OpenStudy (kittiwitti1):

What he said, yes. If not I believe you have to isolate the A.

OpenStudy (littlebird):

I think so, the class is physics.

OpenStudy (kittiwitti1):

Hmm, I wouldn't know then; my Physics teacher said the only necessary prerequisite for Physics is Pre-Calculus

OpenStudy (solomonzelman):

Well, you can kind of make sense here, as of: \(\color{#000000}{ \displaystyle x(t)=At^2 + Bt + C\quad \Longleftrightarrow \quad x(t)=\frac{1}{2}at^2+(v_{0x})t +x_0 }\)

OpenStudy (solomonzelman):

Compare the two, and it should clearly follow that: \(\color{#000000}{ \displaystyle At^2 =\frac{1}{2}at^2 }\)

OpenStudy (solomonzelman):

where lower case a is acceleration.

OpenStudy (littlebird):

We've already been using concepts from calc. The current subject is acceleration due to gravity on an inclined plane. Solomon's idea might work

OpenStudy (solomonzelman):

You get the same thing either based on using the formula for displacement as a comparison, or if you differentiate twice.

OpenStudy (littlebird):

I just noticed that O.o